OpenStudy (anonymous):

need help Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - y = -11 2x + 3y = 5 -4x - 9y = -19 -10y = -30 4x + 3y = 5 2y = -6 7x - 3y = -11 9x = -6 12x - 3y = -33 14x = -28

OpenStudy (anonymous):

nvm i figured it out

OpenStudy (anonymous):

4x - y = -11 2x + 3y = 5 --->(-2) -------------- 4x - y = -11 -4x - 6y = -10 (result of multiplying by -2) -------------add -7y = -21 y = 3 now sub 3 in for y in either equation to find x 2x + 3y = 5 2x + 3(3) = 5 2x + 9 = 5 2x = - 4 x = -2 so your solution is : x = -2 and y = 3 now lets find the one with the same answer.. -10y = -30 y = 3 now sub 3 in for y in the other equation 4x - 9y = -19 4x - 9(3) = -19 4x - 27 = -19 4x = -19 + 27 4x = 8 x = 2 not the first answer choice because the 2 is positive and it is supposed to be negative next one.. 2y = -6 y = -3 not the second answer choice because we got a -3 and its supposed to be positive next one.. 9x = -6 x = - 2/3 not the third answer choice next one.. 14x = -28 x = -2 now sub -2 in for x in the other equation 12x - 3y = -33 12(-2) - 3y = -33 -24 - 3y = -33 -3y = -33 + 24 -3y = - 9 y = 3 It would have to be the last one I tried...lol answer is : the 4th answer choice